# Minds versus Machines: Rethinking Entailment Verification with Language Models

### v1

@@ B.2 Instruction-Finetuned models
These are the more recent “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions. Following are models included in our evaluation.
@@
Chung et al. (2022) It is instruction-tuned from T5 Raffel et al. (2020) on 1.8K+ tasks. We adopt a publicly available version on HuggingFace with model card name “google/flan-t5-xxl”. Flan-T5-xxl is also exposed to data leakage issues. BoolQ Clark et al. (2019), ECQA Aggarwal et al. (2021) and ANLI Nie et al. (2020) have appeared in its training data. However, this is not a serious problem in the finetuning stage because we transform original datasets into entailment verification format before using them for model finetuning.
We extract labels from the model by focusing on the output probabilities of two words, “Yes” and “No”. After applying the softmax function to those two probabilities, we finalize the label as the word with a probability larger than a given threshold.
@@
Brown et al. (2020) It is a general-purpose autoregressive decoder-only LMs accessible via the OpenAI Completions API. [^5] https://platform.openai.com/docs/api-reference/completions We utilize “text-davinci-003” in OpenAI’s API for evaluation, and the label determination procedure is quite similar to the one in Flan-T5-xxl. We monitor the probability of “Yes” and “No” tokens. After applying the softmax function to them, we select the token that attains a probability larger than a threshold to be our label.
@@
OpenAI (2023) It is the latest generative model published by OpenAI which is optimized for creativity and long context inputs. It is accessible via the OpenAI Chat API. [^6] https://platform.openai.com/docs/api-reference/chat We adopt plain “gpt-4” in OpenAI’s API for our experiments. Unlike other models, the output probabilities are not accessible. Alternatively, we firstly constrain the model to predict “Yes” or “No” only by putting some instructions into the prompt. Then, we check the output text. In the ideal case, we can directly use the output text as the label since only “Yes” or “No” is produced. Otherwise, we need to randomly choose one as the label.
B.2.1 Prompts for Robustness Assessment
To assess the robustness of the model in section 2.2, we design different prompt formats but hold the order of premise and hypothesis in the prompt unchanged. Table LABEL:tab:prompt_sensitivity presents all prompts we tested with Flan-T5-xxl and their corresponding averaged results across datasets. The results suggest that the model is insensitive to the variation in the prompt and yields relatively consistent results. This characteristic is maintained in Flan-T5-xxl + Class and Flan-T5-xxl + Rank as well since they are generally constructed based on Flan-T5-xxl.
@@ Appendix C Majority Prediction and Label Imbalance
In Table LABEL:tab:majority_results, we show the performance of an oracle model that predicts the most frequent label in a dataset. For a label-balanced dataset, the macro-F1 score of such a majority prediction model would be 0.67 (precision 0.5 and recall 1.0). The datasets in the evaluation set have some label imbalance, as evidenced by the lower majority label prediction scores. Since we convert existing 3-class NLI and multi-choice QA datasets into our binary classification task format, it inherently has more not support labels. We have more support instances for the rationale datasets since the dataset creators usually only annotate the rationale for the right choice. Specifically, the ECQA dataset only has positive instances, leading to a 1.0 macro-F1 score for majority prediction (support label). Since it has all support labels, any model predicting even a single non support label gets penalized severely, as is seen in ECQA results. Because of this label imbalance in the datasets, we report the macro-F1 scores instead of accuracy or micro-F1.
Appendix D Results on Hypothesis Groundedness
We define the groundedness of a hypothesis based on the type of knowledge required to establish its validity. Using this categorization, we classify the datasets into two categories as follows:
  • Knowledge-based: A knowledge-based hypothesis typically contains entity-grounded or commonsense knowledge. Such a hypothesis is usually well-defined (i.e., the premise is not required to ground it), and models can potentially leverage their internal knowledge to verify the entailment without using the premise. The datasets that contain such knowledge-based hypotheses are FEVER, ANLI, BoolQ, Entailer, and ECQA.
  • Contextual: A contextual hypothesis requires localized knowledge, i.e., the premise/context grounds the hypothesis. These hypotheses cannot be verified on their own without using the premise. Datasets with contextual hypotheses are WaNLI, CosQA, SIQA, DREAM, and RACE.
@@ Appendix D Results on Hypothesis Groundedness
[sections/figures.tex]groundedness
Performance w.r.t. groundedness
Figure LABEL:fig:hypothesis_groundedness shows the aggregate performance across the two hypothesis categories described above. First, we observe that performance on the knowledge-based hypotheses is higher on average than contextual hypotheses for all models (except GPT-3.5 and GPT-4, for which they are comparable). This is likely because models can potentially leverage their internal knowledge stored in the parameters to check the validity of the knowledge-based hypotheses. Such a scenario is harder for contextual cases since the hypothesis solely depends on the context and would require a strong context understanding to predict the hypothesis entailment.
The models in the plot are arranged with increasing complexity, i.e., along the x-axis, either the model size or the amount/variety of pretraining data increases. We observe that the performance on contextual hypotheses almost consistently improves with scaling, whereas the trend is not so strong for the knowledge-based case. This shows that scaling is useful for training LLMs in understanding contexts. However, this does not necessarily improve LLMs’ knowledge retention or retrieval skills.
@@
Appendix E Human Alignment Experiments
@@ E.1 Human Evaluation Details
In each HIT, the annotator was presented with a format exactly like Figure LABEL:fig:eval_annotate, including detailed task descriptions and label explanations. Annotators were expected to read the premise and claim first, then determine the supportiveness of the claim based on the premise and choose the corresponding label. Initially, we only provided three labels — “support”, “irrelevant” and “contradict”. But later, we realized that annotators could not explicitly identify labels for some ambiguous instances where the premise only partially supported or contradicted the claim. Hence, we introduced two weak labels (“partially support” and “partially contradict”) to remedy this issue. When collating results for analysis, we internally combined “support” and “partially support” to be “support”, and the rest to be “not support”, aligning to the standard entailment verification (EV) setup. Each instance was annotated by 3 MTurk annotators, and a majority verdict determined the label. The Fleiss’s kappa score Fleiss (1971) we got was 0.6, indicating a moderate level of agreement among annotators.
@@ E.2 Reasoning Type Annotation Details
In every HIT, we used Figure LABEL:fig:reason_annotate as the reasoning type annotation format. It was an intense job with two tasks involved. The instructions and label explanations were explicitly stated at the beginning of the format. For task 1, after reading the premise and the claim, annotators should decide whether the supportiveness of the claim could be decided by only referring to the information in the premise. If yes, annotators needed to choose the corresponding difficulty level of reasoning about the supportiveness of the claim in task 2. Otherwise, the type of missing information should be decided in task 1, and task 2 did not apply to these cases. As stated in 2.5, we aggregated answers from two tasks and categorized them into four types. Both R1 and R2 were types where the premise contained all necessary information. The difference was that the difficulty level of reasoning for R1 was Easy while for R2 was Moderate. We combined Missing Entity-grounded Information and Missing Commonsense Information to be R3. Finally, the Missing Localized Information label corresponded to R4.
Given that some NLP background knowledge was required to understand task descriptions and label explanations, we recruited Computer Science graduate students instead of the general public to finish this MTurk annotation. Similarly, every instance was assigned to 2 students, and the majority vote determined the label. The Fleiss’s kappa score for this job was 0.62, showing a substantial inter-annotator agreement.
@@ E.3 Reasoning Type Examples
However, unlike R1, the third example of R2 requires three steps of reasoning that “De Baandert was a multi-use stadium.”, “It was mostly used for football matches.” and “The stadium was able to hold 22,000 people.”. The hypothesis can be disproved with those steps because “22,000 people” is just the maximum capacity.
@@ F.1 Training Dataset Selection
To train the Flan-T5-xxl model, we create a training dataset using representative datasets from each category. We pick the ANLI, RACE, and ECQA datasets to represent NLI, contextual QA, and rationale categories, respectively. We note that the amount of training data is quite low for the Rationale category. We also include the StrategyQA Geva et al. (2021) dataset in the training set to alleviate this. Similar to BoolQ, StrategyQA has a Yes/No type of questions and their corresponding explanations. We convert the question-answer pair into a hypothesis using the QA-to-statement converter Chen et al. (2021) as described in Section 2.1.
@@ F.3 Hyperparameters and other details
During training, we select the learning rate from the set $\{7e^{-5},1e^{-4},2e^{-4}\}$, per GPU batch size from the set ${6,8}$, margin $m$ in Equation 2 from the set $\{0.2,0.3,0.5\}$, and warmup ratio $0.1$. The model is trained for 1400 steps on a cluster of 8 A6000 GPUs. We evaluate the model every 200 steps and save the checkpoint if the model shows improvements on a held-out development set.
@@ Appendix G Chain-of-Thought Filtering
G.1 Filtering Examples
Table LABEL:tab:filter_examples presents three CoT reasoning examples, each including two outputs that are kept and three that are filtered out by ranking. According to the table, outputs supported by strong rationales are ranked highly and kept. On the other hand, if the rationale is irrelevant to the prediction (like rationale 3 in example 1), the rationale itself is incomplete (like rationale 5 in example 1), or the rationale supports another option rather than the prediction (like rationale 4 in the example 1), then such output will have a low entailment score leading to a lower ranking and getting filtered out.

### v3

@@ B.2 Instruction-Finetuned models
These are the more recent “general-purpose” language models trained on a collection of NLP tasks described using instructions, leading to generalization abilities to solve unseen tasks described using new instructions. The models included in our evaluation from this category are described below.
@@
Chung et al. (2022) It is instruction-tuned from T5 Raffel et al. (2020) on 1.8K+ tasks. We adopt a publicly available version on HuggingFace Wolf et al. (2020a) with model card name “google/flan-t5-xxl”. Flan-T5-xxl is also exposed to data leakage issues. BoolQ Clark et al. (2019), ECQA Aggarwal et al. (2021), and ANLI Nie et al. (2020) have appeared in its training data. However, this is not a serious problem in the finetuning stage because we transform original datasets into entailment verification format before using them for model finetuning.
We extract labels from the model by focusing on the output probabilities of two words, “Yes” and “No”. After applying the softmax function to those two probabilities, we finalize the label as the word with a probability larger than a given threshold, as described in Equation 1.
@@
Brown et al. (2020) It is a general-purpose autoregressive decoder-only LLM accessible via the OpenAI Completions API. [^9] https://platform.openai.com/docs/api-reference/completions We utilize “text-davinci-003” in OpenAI’s API for evaluation, and the label determination procedure is similar to the one in Flan-T5-xxl.
@@
OpenAI (2023) It is the latest generative model published by OpenAI, which is optimized for creativity and long context inputs. It is accessible via the OpenAI Chat API. [^10] https://platform.openai.com/docs/api-reference/chat We adopt plain “gpt-4” in OpenAI’s API for our experiments. Unlike other models, the output probabilities are not accessible. Thus, we constrain the model to predict a single token as its generated prediction. In the ideal case, we can directly use the output text as the label if “Yes” or “No” is produced. If some other token is generated, we choose one of the labels at random. We note that this occurrence is rare in our GPT-4 evaluation runs.
B.2.1 Prompt Robustness Evaluation
To assess the robustness of the model in section 2.3, we design different prompt formats but hold the order of premise and hypothesis in the prompt unchanged. Table LABEL:tab:prompt_sensitivity presents all prompts we tested with Flan-T5-xxl and their corresponding averaged results across datasets. The results suggest that the Flan-T5-xxl model is robust to the variation in the prompt and yields relatively consistent results. This characteristic is maintained in Flan-T5-xxl + Class and Flan-T5-xxl + Rank as well since they are developed over based on Flan-T5-xxl.
@@ Appendix C Majority Prediction and Label Imbalance
In Table LABEL:tab:majority_results, we show the performance of an oracle model that predicts the most frequent label in a dataset. For a label-balanced dataset, the macro-F1 score of such a majority prediction model would be 0.67 (precision 0.5 and recall 1.0). The datasets in the evaluation set have some label imbalance, as evidenced by the lower majority label prediction scores. Since we convert existing 3-class NLI and multi-choice QA datasets into our binary classification task format, it inherently has more not support labels. We have more support instances for the rationale datasets since the dataset creators usually only annotate the rationale for the right choice. Specifically, the ECQA dataset only has positive instances, leading to a 1.0 macro-F1 score for majority prediction (support label). Since it has all support labels, any model predicting even a single non support label gets penalized severely, as is seen in ECQA results. Because of this label imbalance in the datasets, we report the macro-F1 scores instead of accuracy or micro-F1 metrics.
Appendix D Debatable Cases in Contextual QA
We include two examples of SIQA dataset in Table LABEL:tab:contextual_examples to illustrate the debatable hypotheses in Contextual QA datasets. These examples demonstrate that some less probable hypotheses can also be considered plausible in certain situations. In the first example, option “run screaming” best describes the audience’s reactions after listening to a scary ghost story. However, there are also situations where people are so terrified and start to cry, making the option “cry” somehow debatable. Similarly, in the second example, it is common sense that people will express gratefulness after someone else builds a house for them. But it is reasonable to say that Quinn is intrigued by the house after seeing the amazing architecture. In both cases, apart from the most adequate one, there are still options leading to some partially supported hypotheses, which explains the low human annotation agreement in these cases.
@@ Appendix D Debatable Cases in Contextual QA
[sections/tables.tex]contextualExamples
@@
Appendix E Human Evaluation and Analysis
@@ E.1 Human Evaluation Details
In each HIT, the annotator was presented with a format exactly like Figure LABEL:fig:eval_annotate, including detailed task descriptions and label explanations. Annotators were expected to read the premise and claim first, then determine the supportiveness of the claim based on the premise and choose the corresponding label. Initially, we only provided three labels — “support”, “irrelevant” and “contradict”. But later, we realized that annotators could not explicitly identify labels for some ambiguous instances where the premise only partially supported or contradicted the claim. Hence, we introduced two weak labels (“partially support” and “partially contradict”) to remedy this issue. When collating results for analysis, we internally combined “support” and “partially support” to be “support”, and the rest to be “not support”, aligning to the standard entailment verification (EV) setup. Each instance was annotated by 3 MTurk annotators, and a majority verdict determined the label. To ensure annotation quality, we conducted two rounds of qualification for annotators using the finalized template. In the first round, we used ten questions from the datasets and evaluated 400 mTurk annotators. We retained 100 annotators from this batch with an accuracy greater than 70%. Among remained ones, we repeated this process for another ten questions and selected annotators with an accuracy greater than 80%. Finally, we retained 35 annotators who had annotated the human evaluation set. And the Fleiss’s kappa score Fleiss (1971) we got was 0.6, indicating a moderate level of agreement among annotators.
@@ E.2 Reasoning Type Annotation Details
We annotate the reasoning type of the 100 sampled instances for each dataset with absolute macro-F1 difference $>0.1$ in Table LABEL:tab:human_results. We make this choice intending to attribute the largely misaligned datasets (namely, ANLI, CosQA, SIQA, and Entailer) since some random noise in the annotation and sampling process can potentially also cause some misalignment.
In every HIT, we used Figure LABEL:fig:reason_annotate as the reasoning type annotation format. The instructions and label explanations were explicitly stated at the beginning of the HIT. For task 1, after reading the premise and the claim, annotators had to decide whether the supportiveness of the claim could be decided by only referring to the information in the premise. If yes, annotators needed to choose the corresponding difficulty level of reasoning about the supportiveness of the claim in task 2. Otherwise, the type of missing information had to be decided in task 1, and task 2 did not apply to these cases. We aggregated answers from the two tasks and categorized them into four types. Both R1 and R2 were types where the premise contained all necessary information. The difference was that the difficulty level of reasoning for R1 was Easy while for R2 was Moderate. We combined Missing Entity-grounded Information and Missing Commonsense Information to be R3. Finally, the Missing Localized Information label corresponded to R4.
Given that some strong NLP background knowledge was required to understand task descriptions and label explanations, we recruited Computer Science graduate students instead of the general mTurk workers to finish this annotation. Every instance was assigned to 2 students, and the majority vote determined the label. The Fleiss’s kappa score for this job was 0.62, showing a substantial inter-annotator agreement.
@@ E.3 Reasoning Type Examples
However, unlike R1, the third example of R2 requires three steps of reasoning that “De Baandert was a multi-use stadium.”, “It was mostly used for football matches.”, and “The stadium was able to hold 22,000 people.”. The hypothesis can be disproved with those steps because “22,000 people” is just the maximum capacity.
@@ F.1 Training Dataset Selection
To train the Flan-T5-xxl model, we create a training dataset using representative datasets from each category. We pick the ANLI, RACE, and ECQA datasets to represent NLI, contextual QA, and rationale categories, respectively. We select the datasets with diverse entailment challenges and aim to maximize the total training data. We note that the amount of training data is quite low for the rationale category. Thus, we also include the StrategyQA Geva et al. (2021) dataset in the training set to alleviate this. Similar to BoolQ, StrategyQA has a Yes/No type of questions and their corresponding explanations. We convert the question-answer pair into a hypothesis using the QA-to-statement converter Chen et al. (2021) as described in Section 2.1.
@@ F.3 Hyperparameters and other details
During training, we select the learning rate from the set $\{7e^{-5},1e^{-4},2e^{-4}\}$, per GPU batch size from the set $\{6,8\}$, margin $m$ in Equation 2 from the set $\{0.2,0.3,0.5\}$, and warmup ratio $0.1$. The model is trained for 1400 steps on a cluster of 8 A6000 GPUs. We evaluate the model every 200 steps and save the checkpoint if the model shows improvements on a held-out development set.
@@ Appendix G Chain-of-Thought Filtering
  • +Flan-T5-xxl + Rank + $\mathcal{B}+\textrm{{Flan-T5-xxl + {Rank}}{} + SC}$: This is the same as above, but instead, we use our ranking-finetuned Flan-T5-xxl model for filtering.
@@
[sections/figures.tex]k-selection
G.1 Ablation of Top-k Filtering
We examine our top-$k$ selecting strategy by choosing different values of $k$ and compare their results. We pick up $k$ from $\{3,5,10,20,30\}$ and present the corresponding aggregated results in Figure LABEL:fig:k-selection. The general trend is that, as $k$ increases, the model performance increases first and then starts to decrease, reaching the peak at $k=5$. When $k$ is small, there are only a limited number of CoT rationals, and the majority voting process is vulnerable to potential outliers. On the other hand, if k is too large, then many noisy results are included, leading to poorer performance.
G.2 Examples of Filtered CoTs
Table LABEL:tab:filter_examples presents three CoT reasoning examples, each including two outputs that are kept and three that are filtered out by ranking. According to the table, outputs supported by strong rationales are ranked highly and kept. On the other hand, if the rationale is irrelevant to the prediction (like rationale 3 in example 1), the rationale itself is incomplete (like rationale 5 in example 1), or the rationale supports another option rather than the prediction (like rationale 4 in the example 1), then such output has a low entailment score leading to a lower ranking and getting filtered out.